Kofi and kweku are two brothers

Kofi and kweku are two brothers.kofi is older kweku. given that kofi's age is (5x- 4)years and kweku's age is (2x+1)years,
a.write down an expression, in terms of x,for how much older kofi is than kweku.simplify your answer
b.if kofi is ten years oder than kweku,calculate,
i.the value of x
ii.the ages of kofi and kweku
Answer #1
kofi is older kweku. given that kofi's age is (5x- 4)years and kweku's age is (2x+1)years, a.write down an expression, in terms of x,for how much older kofi is than kweku.simplify your answer (5x-4)-(2x+1) = 3x-5 ----------------------------------------------- b.if kofi is ten years older than kweku,calculate, i.the value of x 3x-3=10 3x=13 x=13/3 ---------------------------------- ii.the ages of kofi and kweku Kofi: 5x-4= 5(13/3)-4 = (65/3)-(12/3) = 53/3 Kweku: 2x+1 = 2(13/3)+1 = 26/3 + 3/3 = 29/3 ============================== Cheers, Stan H.
answered by: Hosiel
Answer #2
(5x-4)-(2x+2) 5x-4-2x+1 5x-2x-4+1 3x-(-3) 3x+3 X=1
source: Knowledge from apizo
answered by: Jessica Henewaa yeboah